### v1

# Describing Globally Distributed Software Architectures for Tax Compliance

Abstract
The company-internal reuse of software components owned by organizational units in different countries is taxable. To comply with international taxation standards, multinational enterprises need to consider a geographical perspective on their software architecture. However, there is no viewpoint that frames the concerns of tax authorities as stakeholders towards a globally distributed software architecture.
In this article, we introduce the reader to the concerns of tax authorities as stakeholders and we investigate how software companies can describe their globally distributed software architectures to tax authorities.
In an in-virtuo experiment, we (1) develop a viewpoint that frames the concerns of tax authorities, (2) create a view of a large-scale, globally distributed microservice architecture from a multinational enterprise, and (3) evaluate the resulting software architecture description with a panel of four tax experts.
The panel of tax experts found that our proposed architectural viewpoint properly and sufficiently frames the concerns of taxation stakeholders. However, the resulting view falls short: Although the architecture description reveals that almost 70% of all reuse relationships between the 2560 microservices in our case are cross-border and, therefore, taxable, unclear jurisdictions of owners and a potentially insufficient definition of ownership introduce significant noise to the view that limits its usefulness and explanatory power of our software architecture description.
Although our software architecture description provides a solid foundation for the subsequent step, namely valuing software components, we stumbled over several theoretical and practical problems when identifying and defining ownership in distributed teams, which requires further interdisciplinary research.
Index Terms:
architecture description, architecture viewpoint, architecture view, software component, microservice, taxation
I Introduction
Code is not tied to a physical, geographical location due to its intangible nature. The code owner of a software component, however, is. Code ownership refers to the concept of assigning responsibility and accountability for a particular piece of code or software component to an individual, a team, or an organizational unit within an organization.
The geographical location of those software component owners has a legal implication that is too often underestimated: Using software components with owners in different countries is taxable as it constitutes an often implicit form of licensing [1]. To comply with the international standard of taxation [2], multinational enterprises must be able to report a geographical perspective on their software architecture. However, there is no architectural viewpoint that frames the concerns of tax authorities as stakeholders towards a globally distributed software architectures which could serve as foundation for a software architecture description.
In this article, we investigate the concerns of tax authorities as stakeholders of globally distributed software architectures and seek an answer to the research question
How can multinational enterprises describe their globally distributed software architectures to tax authorities?
We address this research question in an in-virtuo experiment with a multinational enterprise aiming to report an architecture description of its globally distributed software architecture. Instead of all involved tax authorities, we build a panel of four tax experts that evaluates the usefulness and completeness of the two core elements of the software architecture description, namely the architecture viewpoint and the architecture view. This experimental design enables us to focus on evaluating emerging technical, procedural, and legal issues in practice while reporting software architecture descriptions to tax authorities without any potential legal implications and consequences—which is crucial in this sensitive context.
The main contributions of this study are the following:
  • An introduction to the concerns of tax authorities as stakeholders of globally distributed software architectures
  • A novel software architecture viewpoint specifying the conventions for constructing and using a view for the concerns of tax authorities towards geographically distributed software architectures
  • An empirical view of an industrial globally distributed microservice architecture using our novel viewpoint
  • An evaluation of the software architecture description for tax compliance by a panel of tax experts
Throughout this paper, we use the terminology and definitions of ISO/IEC/IEEE 42010:2022 [3]. In particular, we distinguish between architecture description, architecture view, and architecture viewpoint: An architecture description includes one or more architecture views. An architecture view (or simply, view) addresses one or more of the concerns held by the system’s stakeholders. An architecture view expresses the architecture of the system of interest in accordance with an architecture viewpoint (or simply, viewpoint). There are two aspects to a viewpoint: the concerns it frames for stakeholders and the conventions it establishes on views. Viewpoint conventions can include languages, notations, model kinds, design rules, and/or modelling methods, analysis techniques, and other operations on views.
The remainder of the article is structured as follows: After establishing the background and related work in Section II and discussing our study design as an experimental simulation in Section III, we define the viewpoint (Section IV-B) and how we use it to generate a view for the concern of tax authorities (Section IV-C). In Section V, we present and discuss the evaluation results for our novel viewpoint (Section V-A and the view (Section IV-C) in detail. We conclude the paper with a summary of our findings and the implications for research and practice prior to outlining future work. Figure 1 provides an overview of our study.
[…]
Figure 1: An overview of the structure of our study as in-virtuo experiment and its two components: (1) creating a software architecture description with a viewpoint derived from the concerns of tax authorities and a view of an empirical software architecture using the viewpoint and (2) evaluating the software architecture description with a panel of tax experts.
II Related work
This article connects taxation with software architecture distributed across multiple countries. Therefore, we provide related work on globally distributed software development, ownership in the context of software engineering and taxation, and the emerging field of taxation in software engineering in the following subsections.
II-A Globally distributed software development
Globally distributed software development, especially in large enterprises, is a common practice. Tough competition and a lack of resources motivated many software companies to look for ways to optimize their costs by establishing their own subsidiaries in lower-cost premises locally and nationally or even in other countries or by teaming up with allies that can help in delivering products more cost-effectively. As a result, distributed working across national boundaries has become widespread [4]. However, cross-border collaborations are far from being challenge-free. They are recognized to be considerably more complex than the most complicated one-site development endeavors due to the significant communication and coordination overhead. The solutions to the problems of global projects have evolved through the years. They could be characterized by a long journey from initial problematic experiences of collaborative work through strictly independent work in distributed sites to the evolved ways of working across borders.
Architectural modularization was one of the most prominent early solutions to the coordination and communication challenges in global software engineering. In 1968, Conway published his observation that software architecture cannot be separated from the structure of the organization that develops the software, known as Conway’s law [5]. Inspired by this observation, architectural modularization has emerged as a way to minimize dependencies, structure and allocate tasks, and encapsulate the work in a single geographic development site as much as possible so that communication, coordination, and synchronization needs are minimized [6, 7, 8]. Although the modularity approach to coordination has proven very useful, it has many limitations [9, 10, 11]. While some researchers advocate that strict modularization of work is the possibly best way to alleviate the challenges of distributed work, others argue that work in isolation has many disadvantages [9], such as redundant work, cheap and dirty architectural decisions, misplaced functionality [12, 10] and integration problems [13, 14]. Software components are never truly independent and even the best designs are never error-free, while changes are never completely predictable [14].
Further, strict code ownership is also found to introduce communication overhead and delays [11]. For these reasons, and due to the increased use of agile methods, there has been an increasing interest in collaboratively developing software products. This includes a shift from modularization based on geographically distributed locations towards componentization through microservices [15] on a team level, as well as a trend towards establishing collective ownership of source code and other documents to allow everyone in the project to work on any model or artifact [11]. Collaborative work includes keeping track of which team owns and does what (e.g., through Scrum of Scrum fora or organizing cross-unit demos), submitting contributions to other’s components and reviewing each other’s code [1, 16], and fostering reuse of artifacts to reduce redundant work.
II-B Ownership in software engineering and taxation
Ownership plays a crucial role in taxation in general. However, as we learned in the course of this study (see Section V), the definition of ownership in the context of software engineering and the definition of ownership in taxation overlap but are not congruent due to the different context and perspective towards ownership.
Ownership in software engineering, on the one hand, refers to the assignment of responsibility and accountability for specific portions of a software to individuals or teams within a development organization. There are different perspectives on ownership and autonomy in software development organizations. Software development organizations usually adopt the weak ownership principle [17]. This principle entails assigning ownership of specific components (or microservices) to individual teams. In an ideal scenario, the designated team assumes responsibility as the primary (or exclusive) contributor to the associated component. Nevertheless, empirical analysis of real-world instances often reveals a departure from this idealized model [18, 19, 20, 21]. In practice, a given component may offer support to multiple business streams, undergo development by distinct teams, and necessitate the involvement of specialized teams to address non-functional requirements, such as security and reliability. Consequently, while a specific team retains responsibility for the component, the extent of its contribution to the component can exhibit considerable variability [22].
Tax professionals, on the other hand, begin from a civil law perspective and understand ownership in terms of the legal or economic owner regarding the right of exploitation [2]. While the identification of the legal or economic owner is relevant for balance sheet allocations, in the context of transfer pricing, the legal owner is considered (only) as the starting point for a functional and risk analysis that aims for a profit allocation commensurate with the value-adding activities performed by the legal entities within the group, the so-called functional ownership [2].
Describing and discussing the intersection as well as the difference of definitions is beyond the scope of this work and requires future interdisciplinary research.
II-C Taxation
Taxation in software industry has been debated for many decades [23]. The problem with taxing the final result of software engineering, the software product or service, for example, has shown to be challenging to tackle and is still subject to ongoing and broad discussion [24]. This debate has also reached software engineering, the construction of the software product; at first in the context of inner source, use of open source software development practices and the establishment of open source-like communities within an organization [25]. To address the taxation challenges in inner source software development, Buchner and Riehle proposed a cost-based approach [26]. However, Treidler et al. showed that a cost-based approach is not applicable since inner source does not reflect the function and risk profile, the starting point for any transfer pricing discussion, for a cost-based approach [27].
In a previous article [1], we identified the use of software components owned by teams or individuals who represent separate geographically distributed subsidiaries of the same enterprise as type of licensing—a key intangibles in software engineering that are taxable. For an in-depth discussion on the concerns of tax authorities on software architecture, we refer the reader to Section IV-A in which we discuss a simplified example of a multinational enterprise developing software and how tax authorities become a stakeholder to a globally distributed software architecture.
III Study design
We have designed our study as an in-virtuo experiment (or experimental simulation [28]). In such a setup, the behaviors of actors that a researcher aims to observe and measure are natural and rooted in their professional experience and expertise, but the setting in which these behaviors occur has been specifically created for the purposes of the study. That is, without the study, the setting would not exist. Thus, the term “simulation” refers in this context to the artificially created setting that aims to recreate a concrete type of setting in which the “experiment” (i.e., observed behavior) takes place [28].
In our study, we simulate how software companies can describe their globally distributed software architectures for tax compliance. The simulation consists of two parts: (1) creating a view using our viewpoint at a multinational enterprise and (2) the evaluation by a panel of tax experts in which we observe and measure the usefulness and completeness of an architecture description for tax compliance.
On the one hand, the behaviors of actors we aim to observe and measure are all natural: (1) the multinational enterprise fulfilling its reporting obligations regarding their globally distributed software architecture to comply with international taxation standards, (2) the tax advisors representing and consulting companies and creating a line of reasoning, and (3) tax auditors from tax authorities scrutinizing the line of reasoning. We instructed all participants to imagine it to be in the context of a tax audit.
On the other hand, the setting is specifically created for the purpose of this study, also as a means to ensure the anonymity. Obviously, the evaluation of our proposed architecture description in a real tax audit (i.e., a natural, real setting) is not possible: Tax audits are highly sensitive and, if not brought to court, confidential. Our setting must therefore provide a safe space for all participants without any legal implications for the company or the tax authorities.
However, our setting is not plucked out of thin air but emulates a tax audit as it involves all relevant parties: The panel of four tax experts consists of two tax auditors and two tax advisors representing both opposing perspectives, the company perspective as the provider and the tax authority perspective as the recipient of such documentation. Their expertise and experiences will give us a first evaluation of our preparatory steps and the resulting software architecture description of globally distributed software architectures for tax compliance .
Although this evaluation is a key pillar of our study, our study goes beyond a traditional judgment study as it also describes the practical issues for creating and using a view based on our novel viewpoint in an empirical case. We discuss our approach to the limitations in Section VII.
Last but not least, we would like to emphasize that an in-depth legal discussion and evaluation of our work is beyond our reach.
III-A Data collection
All involved parties want to remain anonymous including our case company used as a source for creating our view. Therefore, we are not able to describe the case and the data collection in detail. However, we may well argue that our case company is exemplary for a multinational enterprise developing software.
In brief, for creating the view, we extracted the software architecture and the team information from an internal system that extracts the component structure of the system from the code repositories on a daily basis.
For the evaluation of the usefulness and completeness of our software architecture description for its intended purpose, we conducted semi-structured interviews with four tax experts. The main goal of the evaluation is to seek generalizability over the responses rather than generalizability to a population of actors [28]. We ensured privacy and confidentiality for all actors (i.e., the involved companies, tax authorities, and tax advisors) by conducting the interviews separately and removing all company-specific information. We also altered the geographical distribution of the subsidiaries to make our company not identifiable in any sense on the one hand but also to give the reader a better impression of globally distributed software architecture on the other hand.
The interview guidelines are publicly available [^1] https://zenodo.org/records/10214942. The group of four tax experts represent the two opposing parties in a tax audit: internal or external tax advisors representing the company and tax auditors representing the tax authorities.
We conducted two pilot interviews to address didactical challenges since the interviews were conducted with tax experts with no software engineering background. A pair of a co-author with a taxation background and a co-author with a software engineering background conducted the interviews with the tax experts virtually between 2023-10-24 and 2023-11-03. Table I provides an overview of our interviewees and their background. We recorded all interview sessions.
TABLE I: An overview of the interviewees and their backgrounds.
ID Role Background
A Tax auditor
Former tax auditor, majoring in international taxation, OECD advisor for transfer pricing from 2016 to 2018, currently not involved in transfer pricing
K Tax advisor
Transfer pricing advisor, previously transfer pricing expert at one of the Big Four accounting firms for 20 years
M Tax auditor
Tax auditor and key expert on transfer pricing for other tax auditors for more than eight years
R Tax advisor
Managing director at a transfer price consulting firm; tax inspector before, 35 years of experience in international taxation
III-B Data analysis
As our interview guidelines show, large parts of the interviews were dedicated to providing contextual software engineering understanding for the tax experts. Therefore, we translated and paraphrased only the answers from the interviewees. While one researcher of the interview team analyzed the recorded interviews using thematic analysis [29], the second one ensured the quality of the findings and mappings. After the coding, we again reflected on the findings.
IV A software architecture description for tax compliance
In the following subsections, we provide a gentle introduction to international taxation for software architects and specify the set of conventions for the creation, interpretation, and use of an architecture view to frame the concern of tax authorities.
IV-A International taxation standards as concern and tax authorities as stakeholder
Consider the fictional devnullsoft Group, a multinational enterprise that develops and sells a software product. The software is split up into 18 components. Each component has an owning team that is responsible and accountable for its components. The teams are assigned to one of the legal entities of the enterprise in three countries depending on the location of the team members: devnullsoft GmbH in Germany, devnullsoft AB in Sweden, and devnullsoft Ltd. in the UK. In Figure 2, we provide a schematic overview of a simplified globally distributed software architecture with owning teams assigned to three different subsidiaries in Germany, Sweden, and the UK. Using code components across different jurisdictions is taxable.
[…]
Figure 2: A schematic overview of a globally distributed software architecture developed by the fictional devnullsoft Group, a multinational software enterprise with three subsidiaries developing software. Using code components across different jurisdictions is taxable.
Only the German devnullsoft GmbH sells this resulting product to customers. Without any further consideration, solely the German devnullsoft GmbH generates profits, which are then fully taxed in Germany according to German law. The Swedish and English tax authorities are left out in the cold because devnullsoft AB and devnullsoft Ltd. have no share of the profit that could be taxed in Sweden or in the UK, although both subsidiaries maintain, test, and develop, or simply own components that are crucial for the software product that made the success of the software possible.
To avoid this scenario and to provide a common ground for international taxation, reducing uncertainty for multinational enterprises, and preventing tax avoidance through profit shifting, nearly all countries in the world agreed on and implemented the so-called arm’s length principle as defined in the OECD Transfer Pricing Guidelines for Multinational Enterprises and Tax Administrations [2]. This has become the guiding principle and the de-facto standard for the taxation of multinational enterprises that requires associated enterprises to operate as if not associated and regular participants in the market from a taxation perspective. To comply with the arm’s length principle, multinational enterprises must pay a so-called transfer price. Transfer prices are the prices at which an enterprise transfers physical goods or—for software much more relevant—intangibles or provides services to associated enterprises. Those transfer prices are established on a market value basis and aim to avoid profit shifts from high to low tax regions. Although other types of intangibles, like cross-border code contributions, code reviews, bug reports, etc., are also relevant from a taxation point of view, we focus in this article on another highly relevant type of intangible in software engineering: licenses.
Although maybe not even explicit, the company-internal use and reuse of software components is, in fact, an instance of licensing. A software license is a legal instrument governing the use or redistribution of software. Software licenses and their obligations are widely discussed in the context of open-source software. But in contrast to open source, components and their use are protected by the confines of a company. Third parties are excluded from any use or redistribution right of the software components. Without further considerations—i.e., transfer prices—this exclusion contradicts the arm’s length principle and makes multinational enterprises like our devnullsoft Group not compliant.
In this way, tax authorities become a stakeholder in globally distributed software architectures as they have the concern of identifying cross-border reuse of software components as an instance of taxable licensing. Determining the actual transfer pricing for the cross-border reuse of software components is a vast and separate challenge on its own and not part of this work. Our work focuses on the architectural description comprising all necessary information for reporting the software architecture to tax authorities.
To conclude, the concern of tax authorities on software architecture is the (potentially implicit) license structure of the software-intensive product within the multinational enterprise. From this concern, we can derive three main questions:
  • How is the software-intensive product structured?
  • What legal entities (implicitly) license those substructures of the software?
  • Where are those legal entities geographically located?
All three questions must be answered in the context of software architecture.
IV-B Architecture viewpoint
To address the concern of tax authorities and to answer the three derived questions, our viewpoint requires four elements: Defining the component and component owner, decomposing the component structure of the software product, identifying component owners, and identifying the jurisdictions of the component owners.
IV-B1 Defining component and component owner
A software component is a self-contained, reusable piece of software that encapsulates the internal construction and exposes its functionality through a well-defined interface so other components can use the functionality. That enables software components to be used, developed, tested, and maintained individually by different teams and separately from the rest of the software product. Software components can take many forms, including libraries, modules, classes, functions, or even entire applications.
Component ownership refers to the concept of assigning responsibility and accountability for a particular software component to an individual or an organizational unit within an organization. Although there are different notions of code ownership (for example, strong, weak, or collective code ownership [17]), component ownership is more strictly defined in this context: We requires only one organizational unit is ultimately responsible and accountable for a software component.
The definition of a component or a component owner is company-specific and may vary. Thus, an in-depth rationale for the definitions in the context of the reporting company is required.
IV-B2 Decomposing the component structure of the software product
To reveal the component structure and the dependencies of software components, the software product must be decomposed into the component structure. This results in a dependency graph reflecting the dependencies between the software components. The type of dependency depends on the definition of a software component. Therefore, a rationale is required.
IV-B3 Identifying component owners
Each component must be assigned to one owner. This usually results in a table of components and their owners. As discussed previously, the mapping of owner to component is a one-to-many mapping.
IV-B4 Identifying jurisdictions of component owners
Each organizational unit (or individual) owning a software component must be assigned to one jurisdiction. In most cases, a jurisdiction is the country where the local subsidiary is located. This usually results in a table of component owners and countries. As discussed previously, mapping an owner to a country is one-to-one.
IV-C Architecture view
In this section, we discuss the practical view construction using the presented viewpoint at a large multinational enterprise.
In detail, we use microservices as the most suitable abstraction layer for software components for the specific software architecture in our case. A microservice represents a discrete piece of a larger application’s functionality and operates as a self-contained unit. Microservices are typically designed to be lightweight, modular, and independently scalable, allowing for greater flexibility and agility in software development [15]. Therefore, microservices meet the definition of a software component.
The software architecture of the software product consists of 2560 different microservices. We consider only microservices in production; experimental or other non-production microservices are excluded. Microservices in production imply that those microservices face the customer directly or indirectly.
In our case company, a component owner is responsible and accountable for its software component. That implies ownership includes functional and non-functional requirements, such as security, reliability, and quality, and extends throughout the software lifecycle, including development, testing, and maintenance. It also involves all aspects of operating that code in production, such as deployment, monitoring, troubleshooting, and remediation of operational and security incidents.
In our case company, each microservice is usually owned by a single team. We excluded 42 microservices (1.64%), each owned by one individual because the locations of individuals were not accessible for this research due to privacy concerns. Those remaining 2518 microservices have 16533 dependencies between them and are owned by 336 different teams.
The geographical locations of the teams are self-assigned: Teams report their location in a central system. We find this self-reporting most realistic in the world of distributed and remote teams. However, we were not able to retrieve or confirm the actual geographical location of 140 teams (41.16%) with this self-reporting approach. This is mostly due to remote and distributed teams that do not have a single location that we could assign to a single jurisdiction. This affected 8097 use relationships, i.e., 48.94% of all use relationships.
Due to the large amount of data, we present the architecture view in two aggregated visualizations instead of a plain list, as graph and as table.
The directed graph as in Figure 3 depicts jurisdictions (countries) as vertices and dependencies as directed edges. An edge goes from the component owner to the component user. We aggregate the number of dependencies between countries as edge weights.
[…]
Figure 3: The geographical distribution of the microservice architecture of our case company using our architectural viewpoint. The arrows indicate the direction of the dependency from microservice user to the microservice owner. The countries are partially substituted by equivalent countries to ensure confidentiality for the company.
This graph, however, lacks of the information of unclear locations. Therefore, we provide also a tabular form of the graph including the information of missing or unclear location. Table II depicts this information.
TABLE II: An equivalent tabular representation of Figure 3 but including unknown dependencies.
Component owner
DEU GBR NLD FRA USA N/A
Component user DEU 2 2 0 0 0 4
GBR 15 164 2 261 43 141
NLD 3 6 19 11 5 8
FRA 24 108 21 4069 850 1767
USA 14 24 15 1130 1648 642
N/A 27 70 14 2283 970 2171
V Evaluation results
In this section, we report the evaluation results from the panel of tax experts (see Section III for more details) evaluating the software architecture description with its viewpoint that framed the concerns of tax authorities and the view of software architecture from our industry partner.
V-A Viewpoint
Our panel confirmed the correctness of our viewpoint and agreed with the distilled three questions raised from the concern of international taxation. As interviewee A stated: I think it is possible to reveal the license structure [by your viewpoint], at least by 90%, there will always be fuzziness.
However, the panel also agreed that this viewpoint can only be the beginning. You are approaching it from a taxation perspective from a very detailed level, as interviewee R said. [Your viewpoint] is an important first step towards understanding who is actually contributing to the software product. But to answer the follow-up question, which is the valuation of the software component, you need to add another layer. [Components] are not additive. In the words of interviewee K: Value is not counting components. The classification of software components potentially reflects how easily the components can be replaced or how differentiating the components from competitors are.
In this context, all participants emphasized the importance of ownership from a taxation perspective. Interviewee M and interviewee K highlighted accountability as a key aspect. They raised the question of who is actually accountable for the software product or the components from a legal perspective.
Our panel emphasized that although aggregating over jurisdictions can be acceptable in our case since there is only one subsidiary per jurisdiction, it does not generalize: If there is more than one subsidiary in a jurisdiction, aggregating over countries is not applicable. Interviewee K also mentioned that assigning legal entities to jurisdiction can become complicated referring to so-called permanent establishments, where, in complex situations, subsidiaries may have international branches and sites fall into jurisdiction X but belong to jurisdiction Y. Assigning jurisdiction is complex. There are gray areas. However, those special cases are likely not relevant for those considerations.
Interviewee M wondered how often software architectures may change and commented that the viewpoint does not consider such changes over time but represents a snapshot of software architecture. As a tax auditor, the interviewee would question the representativeness of the presented snapshot.
V-B View
When creating a view using the viewpoint, our panel identified two major issues: (1) the significant amount of unclear links between owner and jurisdiction and (2) the missing qualitative insights into the software architecture.
The first major issue refers to the problem that we were not able to identify the jurisdiction of 41.16% of all owners in our view on the given software architecture. Although the data source, the company-internal platform for teams, is promising (I go along with that one. […] That is, in my opinion, an excellent data source) since the data is not biased from a taxation perspective (interviewee M), the amount of unclear jurisdiction of owners is too high. The experts suggested different additions (jurisdiction of the team manager) or complements (a cascade of different other data sources, for example, a questionnaire sent out to all owners or geographical distribution of the team to approximate the jurisdiction of the owners). Although missing some percentages is no problem, but half of the dependencies missing is too much as interviewee M explicitly said. 50% not assignable implies that 50% of all transfer prices are not properly assigned. […] Without data, we are in vacuo. (interviewee A).
In this context, interviewee M and interviewee A emphasized the importance of traceability in tax audits. Therefore, it would also be better to reveal the process and characteristics of how teams came to their self-assignment. Indications like regular socializing in a certain country can be such an indicator. In this context, the panel agreed that the physical location of the team members and developers can be secondary from a transfer pricing perspective.
Our panel appreciated the graph representation and found it useful and utilizable. For example, interviewee M said, I like the [graph] visualization; it is helpful. Interviewee A detailed that the visualization as a graph is helpful for a high-level view, a good starting point for looking into details.
The tabular representation was seen as a useful extension of the graph. Interviewee M preferred the presented absolute numbers over relative numbers since absolute numbers represent the raw information and relative numbers are easy to calculate anyway.
The second major issue raised by our experts is grounded in the quantitative nature of our visualization as graph and as table. The panel of tax experts unanimously warned that the numbers may be misinterpreted as relative value added to the software product. As already discussed in the previous section, considering all components equal is not an acceptable simplification. The edge weight does not reflect the relevance as stated by interviewee R. It requires a classification of the components (interviewee K).
Besides those two major issues, the panel also raised further issues. Interviewee R wondered about the effort required to generate those insights. The interviewee emphasized that the cost-benefit ratio must be positive. Another important remark was related to our decision to exclude non-production components. Interviewee M reminded us that our ultimate goal is to identify the value of software components. Although not in non-production or completely experimental, those components can have a significant and not negligible value for the company. Therefore, excluding them removes explanatory power from the software architecture description.
V-C Software architecture description
At the end of each interview, we asked the participants about their general and high-level impression of our software architecture description with its viewpoint and view in its current state.
The overall impression was positive and, according to our panel, the software architecture description fulfilled its main purpose at its current state to a large extent. Although transfer pricing is not a science, it is art, interviewee R correctly summarized our intention as follows: This is just fact-driven, it is no interpretation. When you ask people, they may or may not know who is involved. The nice thing about it that it is not a subjective view, it is about facts, it is about technical details. This is what is used if you like it or not. That is relevant because you have objectivized the contributions [to the software product]. Those are the parties involved. This allows you to have a very objective starting point to consider the potential parties you have to consider in greater detail. The visualization raises awareness of parties involved in the software development. The panel also confirmed the importance of software architecture for taxation: I think software architecture is the right perspective to understand how software is composed conceptionally (interviewee K). Interviewee R highlighted the need for software that can manage the complexity and support the transfer pricing processes. Interviewee M also appreciated the company’s efforts in collecting and presenting the data in a tax audit: A company with such detailed and thorough documentation would be among the top 10 or top 5 percent of what a tax auditor usually sees. Interviewee K agrees: I really like the approach aiming for objectiveness. Let’s face it: Sometimes there are preposterous debates in tax audits.
VI Discussion
In this section, we critically reflect on the evaluation results and feedback from our panel of tax experts and what it means for our software architecture description and future work.
We found ambiguous assignments of owners to jurisdiction in about half of all dependencies. We expected that this significant amount of unclear jurisdictions per team create a strong headwind. Statistical estimators cannot be applied to make any assumptions on the jurisdictions because we have a strong bias: The owners with missing or unclear jurisdictions come mostly from teams that have team members from different subsidiaries or are completely virtual. Those are inherently relevant for taxation purposes and need further investigation.
The discussion on ownership, however, cast doubt on our definition of ownership. We learned that the term ownership triggers a certain association in the context of taxation: Tax experts think of legal ownership when they read ownership. Although we emphasized in our viewpoint that we require an owner to be accountable and responsible, the term ownership led to different misunderstandings and triggered in-depth legal discussions that go beyond the scope of this paper. We conclude that our definition of ownership can be indeed too weak: Our understanding of accountability may fall short and does not suffice the legal accountability. Future research needs to address the legal aspects of code (component) ownership.
We were surprised that the quantitative nature of our description was conceived that all components are normalized and of equal importance and relevance for the software product. Interviewee R even suggested removing the numbers completely and just indicating by the edges that there is a dependency that needs further investigation to avoid this misconception. Although we agree that those numbers can misinterpreted in such a way that all components are equally important and relevant, we still believe that the weighted edges provide value for prioritizing: The more dependencies, the more potential uncertainty is in a tax audit that needs to be addressed. To avoid the impression that dependencies are additive, we consider replacing the absolute numbers with intervals (e.g., (0,10)010(0,10)( 0 , 10 ), (10,100)10100(10,100)( 10 , 100 ), and ($100,\infty)( 100$ , ∞ )) in future iterations of the software architecture description.
We are not aware of studies addressing at what rate or to what extent microservices architectures or software architecture in general change over time (a fiscal year), which would also may be not generalizable and context-specific anyway. Since the effort to create the views once the data sources are identified is reasonable, we suggest regularly generating snapshots of the software architecture. Furthermore, we hypothesize that changes in software architecture also may trigger changes in ownership, directly or indirectly.
VII Limitations
In this section, we critically reflect on the two components of our simulation-based approach: the software architecture description from our industry partner and the evaluation of the software architecture description in the contrived setting.
VII-A Software architecture description
We created the view as described in the paper. We relied on historical information extracted from an internal system that extracts the software product’s component structure (which also comprises microservices) from the code repositories. We believe that this system is a reliable and trustworthy data source for extracting the data we need since the data is also used internally for other purposes. The internal team responsible for maintaining the system helped us to avoid misinterpretation of the data. The date of extraction was randomly selected from the first two quarters in 2023. Although we did not collect this data, we assume that the microservice architecture does not change at a daily rate but is stable at least for over weeks.
VII-B Contrived setting
The evaluation of the software architecture description for tax compliance highly depends on our tax panel. The contrived setting and the strict anonymization of the case company required us to remove context and potentially relevant information, which includes the specific business model, the revenue stream, and organizational structure of our case company. Not being able to provide that information limits the generalizability of our evaluation. However, we believe that we provided enough context and information for the panel to provide solid and high-level feedback for the early stages of our research.
Both representatives for the tax authorities come from only one country. Although the taxation standards are international and our panel consists of tax experts advising taxation internationally, our study may not fully capture the nuances in interpreting the standards. Therefore, our results only apply to the fundamentals of describing software architectures to tax authorities. Future research is required to address the nuances and country-specific specialization of international taxation.
To maintain strict anonymity for the company, we decided not to include participants from the company. The author team spoke on their behalf. In future work addressing the subsequent questions of valuing software components, the involvement of internals will become crucial. We believe that in this early stage of the software architecture description, the author team who extracted the data brought enough context to the table.
The interviews were held in English and German. To avoid translation mistakes and misunderstandings, the second participating interviewer checked the findings after coding and the final version of the paper was sent to all interviewees, asking for confirmation.
VIII Conclusion
In this study, we proposed and evaluated a novel software architecture description for reporting globally distributed software architecture for tax compliance. The architecture description of our case company reveals that almost 70% of all reuse relationships between the 2560 microservices in our case are cross-border and, therefore, are taxable.
The panel of tax experts found that our proposed architectural viewpoint properly and sufficiently frames the concerns of taxation stakeholders. The panel also appreciated the evidence-based, empirical approach for visualizing the software architectures to tax experts as software-engineering laymen.
However, we identified several challenges that must be addressed in future work. The most critical and likely most complex issue to solve is our insufficient definition and understanding of ownership, which needs to be rethought for taxation purposes in this line of research. Also the question what does or will constitute to the software product and what not deserves further attention. Last, but not least, the subsequent step, adding value to each component, is yet another unsolved problem in the context of taxing software-intensive products. We believe that the software architecture research community can play a vivid role in addressing those challenges and paving the way for tax compliance in software engineering.
At this point, we would like to emphasize that the quality and rigor of the data curation at our case company is extraordinary and, from our experience, not industry standard. Not always, companies can rely on such a solid foundation grounded in historical and accurate data on their software architectures. This uncertainty introduces further risks and lets tax audits stray away from an empirical-based, data-driven discussion on the value of software. Although this research is only the beginning, we recommend companies start continuously collecting information about their software architecture and ownership. Accurately collecting this information is never trivial but nearly impossible in hindsight for complex software systems.
Although there is no black or white in tax audits but a lot of gray, as interviewee R summed up, we hope to shed some light on the discussions in tax audits and lay the foundation for a data-driven approach for taxing globally distributed software architectures.
Acknowledgment
We would to thank our industry partner and the panel of tax experts for their great support which made this research possible.
This work was supported by the KKS Foundation through the SERT Project (Research Profile Grant 2018/010) at Blekinge Institute of Technology.